OpenStudy (anonymous):

how do I find the inverse of y=(0.5)^(x+2)???

OpenStudy (anonymous):

\[y=(0.5)^{x+2}\] x = (0.5)^(y+2) logx = y+2log0.5 logx/log0.5 = y+2log0.5/log0.5 (logx/log0.5) - 2 = y but the correct answer is \[\sqrt[(x-2)]{0.5}\]

OpenStudy (anonymous):

heelp D:

OpenStudy (anonymous):

@asnaseer do you have a clue as to how to get the correct answer? :$

OpenStudy (asnaseer):

how sure are you that the correct answer is \(\sqrt[(x-2)]{0.5}\)?

OpenStudy (anonymous):

sure as in it's written in the back of my book...

OpenStudy (asnaseer):

well it does not look correct to me.

OpenStudy (asnaseer):

maybe there is a printing error in your book?

OpenStudy (anonymous):

really? oh.. okay, thank you!